Using MATLAB, generate a sequence of L=4000 independent Gaussian random variable with variance σ^2 = N0/2.
MATLAB code to generate one Gaussian random variable with variance σ^2: \sqrt(σ^2)∗randn.
Make sure that N0 has to be set to a proper value, depending on the value Eb/N0 (dB) you’re
working on.
Answer #1

ebno=15; %Eb/No is 15 dB

eb=0.05; % bit energy is set as 50mW

L=4000;

n_var=10^(-ebno/10)*eb; % noise variance caluculated from %Eb/No.

grv1=sqrt(n_var)*randn(L,1); % generating a real Gaussian %random variable with required noise variance.

grv2=sqrt(n_var/2)*(randn(L,1)+sqrt(-1)*randn(L,1)); % %generating a complex gaussian random variable with %required noise variance.